how do you figure out the square root of 2 over the square root of 8? please show all steps
Answer by rwm(914)

sqrt(2)/sqrt(8)
sqrt(8)=sqrt(4*2)=2sqrt(2)
so we have sqrt(2)/2 sqrt(2)=1/2